# KVLink: Accelerating Large Language Models via Efficient KV Cache Reuse

### v1

@@ 3.1 Experiment Setup
3.2 Main Results
We highlight the key observations below. First, KVLink consistently outperforms all baselines across all datasets, demonstrating its strong effectiveness. In QA tasks, our approach surpasses the best baseline, BlockAttention, with up to 6% higher accuracy. For other baselines, the gap is more significant. The performance gap is even larger when compared to other baselines. Notably, our method not only outperforms the original Llama models without fine-tuning but also achieves accuracy close to fine-tuned Llama, which is evaluated with a fully concatenated context.
3.3 Inference Efficiency Evaluation
Figure 4: Inference speed comparison with ten reused contexts of varying lengths. Both KVLink1 and KVLink5 show considerably lower Time-to-First-Token (TTFT) than standard decoding as context size grows, demonstrating the efficiency gains from reusing precomputed KV caches.
A key objective of KVLink is to reduce the computational overhead incurred by repeatedly encoding long contexts. To demonstrate these efficiency gains, we measure the Time-to-First-Token (TTFT) when reusing the precomputed KV cache for ten documents of varying lengths. For a realistic assessment, we include the overhead of loading the precomputed KV caches onto the GPU. Each measurement is averaged over 100 runs, with an initial 10 warm-up trials to eliminate memory allocation overhead, following the setup in  (Xiao et al., 2024).
Figure 4 compares our methods, KVLink1 and KVLink5, against standard decoding, which encodes the entire context for each query. The key observations are as follows. First, reusing the precomputed KV cache significantly improves LLM inference efficiency. To leverage the precomputed KV cache, we perform three main operations: (1) loading KV caches from CPU to GPU, (2) applying new positional encoding, and (3) encoding the link tokens for each document. All of these operations introduce only negligible latency, reducing TTFT by 60%–90% compared to standard decoding.
Second, across all context sizes, both KVLink1 and KVLink5 consistently achieve substantially lower TTFT relative to standard decoding. As the reused context length increases, the TTFT gap further widens, highlighting the scalability of our approach. Notably, when the context length reaches 5,000 tokens, KVLink decreases inference latency by 90%. These results confirm that KVLink effectively mitigates the computational cost of large-scale context encoding and is well-suited for scenarios requiring fast response times.
3.4 General Capability Preservation
While our primary objective is to enable efficient KV cache reuse, we also verify whether these modifications maintain the model’s general reasoning and instruction-following capabilities. Table 2 presents the results for two model variants: Llama-3.2-1B and Llama-3.2-3B. We compare the original models, their fine-tuned counterparts, and our method with different numbers of link tokens (KVLink0, KVLink1, and KVLink5).
@@ 3.4 General Capability Preservation
Second, while there are minor drops in certain benchmarks (e.g., ARC-C and Winogrande) compared to the original models, these differences remain within a reasonable range, typically less than 3%. We expect that performance could be further improved by refining the data mixture and incorporating additional supervised fine-tuning on reasoning and instruction-following tasks, which we leave for future work.
@@ 4 Related Work
Efficient inference for LLMs.
As model sizes grow, serving LLMs efficiently becomes increasingly challenging. Previous work has tackled this via model pruning (Men et al., 2024; Sreenivas et al., 2024; Xia et al., 2024; Hou et al., 2025), quantization (Lin et al., 2024; van Baalen et al., 2024; Frantar et al., 2022; Xiao et al., 2023a; Yao et al., 2022; Dettmers et al., 2022), and optimized decoding algorithms such as non-autoregressive decoding (Santilli et al., 2023), speculative decoding (Kim et al., 2024; Elhoushi et al., 2024; Miao et al., 2023), or early-exiting (He et al., 2021; Kong et al., 2022). Like these approaches, our method also aims to enhance LLM inference efficiency by reducing inference-time computation.
KV cache reuse.
Recent work has explored reusing precomputed KV caches across multiple queries, often focusing on prefix-only reuse (Jin et al., 2024; Liu et al., 2024b; Zheng et al., 2023). A few methods extend reuse to arbitrary positions, aiming for greater efficiency. PromptCache(Gim et al., 2024) allows KV cache reuse with discontinuous position IDs but ignores cross-chunk attention, causing performance drops. It also relies on Prompt Markup Language (PML), limiting flexibility across tasks. CacheBlend reduces positional mismatches by selectively recomputing caches for tokens with the largest value-state variance; however, it suffers from performance degradation and the non-trivial cost of recomputing 10%–20% of tokens (plus all tokens at the first layer). BlockAttention(Sun et al., 2024) removes positional embeddings by re-encoding KV caches and directly fine-tunes on block-attention data, but lacks strategies to better connect reused contexts or refine data processing, leaving performance below ideal levels.
KV cache quantization and eviction.
While our method focuses on accelerating inference by reusing KV caches, a complementary line of work aims to reduce memory overhead through cache eviction and quantization. For instance,  (Zhang et al., 2024a; Oren et al., 2024; Xiao et al., 2023b; Ge et al., 2023) introduces an eviction policy that significantly lowers the memory footprint, while  (Liu et al., 2024d; Xiao et al., 2023a; Lin et al., 2024; Frantar et al., 2022; Kim et al., 2023; Zhao et al., 2024; Sheng et al., 2023) investigate quantizing KV caches to minimize storage requirements with only minor performance loss. Notably, our approach can be seamlessly integrated with these state-of-the-art eviction and quantization strategies, thereby addressing both speed and memory efficiency.
@@ 5 Conclusion
In this paper, we introduced a method to improve LLM efficiency by reusing pre-computed KV caches. With precomputed KV caches for retrieved documents or context segments, our method can avoid redundant computation for overlapping contexts across different queries. In the future, we will refine our data mixture and investigate optimal fine-tuning strategies to further enhance performance. Additionally, we plan to explore real-world deployment scenarios to maximize the inference efficiency of LLMs.
6 Limitations
Although our approach achieves state-of-the-art performance while improving the inference efficiency, it still has several limitations. First, our method involves fine-tuning to adapt LLMs to the new paradigm of separately encoded documents. This could be costly for very large models. Second, our method leverages the pre-computed KV cache of retrieved documents or context segments for efficient inference. When the number of documents or context segments increases, there could be a significant overhead on CPU memory or disk storage. However, as the number of documents or context segments increases, it may introduce significant overhead in CPU memory or disk storage. Efficiently managing the storage and loading costs of the pre-computed KV cache remains a challenge for large-scale deployment.
7 Societal Impact
@@ A.2 Implementation Details
Figure 7 illustrates the system prompts we use for each task category during training. These prompts are generated by GPT-4 (Achiam et al., 2023), and are randomly picked during training. Additionally, for the QA training data, we also shuffle the retrieved reused documents in the context A.5.
@@
For QA evaluation with separately encoded KV cache, we adopt the accuracy as the evaluation metric, following (Sun et al., 2024; Liu et al., 2024a; Kandpal et al., 2023; Mallen et al., 2022), which considers the prediction is correct if any sub-string of the prediction exactly matches any gold answer.
For the evaluation of general capability preservation, we use LM Evaluation Harness (Gao et al., 2024). We use few-shot examples for GSM8K and MMLU(8-shot for GSM8K and 5-shot for MMLU).
@@ A.3 More Implementation Details of Baselines
For BlockAttention, we process the training data with context reuse in 3 by separately encoding each reused document or context segment and concatenating the caches directly without inserting any special tokens in between. For the data with no context reuse, the data processing is the same as KVLink.
@@ A.4 Ablation Studies on Data Mixure
Table 4: Performance on different training data mixture.
@@ A.5 Impact of Answer Document Position
Empirically, we have observed that directly constructing the QA training data using Contriever (Izacard et al., 2021) retrieved relevant documents is not ideal because the retriever typically places the answer-containing document near the front based on the relevance score. As a result, the fine-tuned model “learns” to find answers at the beginning of the context rather than reasoning across the entire reused context.
@@
To validate our statement, we fine-tune the Llama-3.2-3B base model on the same training data without shuffling as BlockAttention and evaluate under the (Liu et al., 2024a) setup. In (Liu et al., 2024a), the test set for NaturalQuestions is also built with Contriever-retrieved contexts but places the document containing the correct answer at varying positions. As shown in Figure 5, the model’s performance drops significantly whenever the ground-truth document is located further back in the context.
A.6 Submission Checklist
@@
Various datasets are included for training and evaluation, and their licenses are listed below. The IFEval, NaturalQuestions, 2WikiMQA, TriviaQA, HotpotQA dataset is released under the Apache License 2.0. HumanEval, MMLU, GSM8K, HellaSwag, WinoGrande, XSum and LM-Evaluation-Harness are under the MIT License. The ARC dataset is provided under the Creative Commons Attribution Share Alike 4.0 license, and SciQ is under the Creative Commons Attribution Non-Commercial 3.0 license. The DaringAnteater dataset is under the license of Creative Commons Attribution 4.0. The MuSiQue dataset is under the Creative Commons Attribution 4.0 International license. The Samsum dataset is under the Creative Commons Attribution Non Commercial No Derivatives 4.0 license. The MultiNews dataset is under a legal agreement with LILY LAB. PiQA is licensed under the Academic Free License v. 3.0. The Tulu3-sft-mixture and fineweb datasets are under the Open Data Commons License Attribution family license.
@@
All datasets and software packages in this study are used strictly for their intended purpose—namely, training and evaluating LLMs. We confirm that no personal identifying information or offensive content appears in our materials. Additionally, we employed GPT-4 to assist with proofreading and improving clarity throughout the text. Nevertheless, the research concepts, analysis, and original writing remain fully authored by us.

### v2

@@ 3.4 General Capability Preservation
Llama-3.1-8B
Original Llama 85.1 68.0 85.1 78.7 81.9 51.7 79.8 96.7 73.7 59.1
Finetuned Llama 75.4 64.3 82.7 75.6 80.9 48.3 80.3 96.3 72.9 58.6
KVLink 0 75.2 64.8 82.8 75.8 82.1 50.8 80.0 96.8 72.5 58.9
KVLink 1 75.5 64.0 82.4 75.4 82.2 51.1 80.7 96.8 72.8 59.0
KVLink 5 75.7 64.0 83.0 76.2 81.5 50.4 80.3 96.6 72.4 59.1
@@
Second, while there are minor drops in certain benchmarks (e.g., ARC-C and Winogrande) compared to the original models, these differences remain within a reasonable range, typically less than 3%. We expect that performance could be further improved by refining the data mixture and incorporating additional data on reasoning and instruction-following tasks, which we leave for future work.
3.5 KVLink with Cache Compression
In this section, we further evaluate the effectiveness of our method when combined with KV cache compression techniques, which help reduce the overhead of storing and loading precomputed KV caches. For both compression methods described in Section 2.4, we experiment with 50% and 75% compression rates. We fine-tune the Llama-3.2-1B model using our method and the baseline models on compressed KV caches using the training sets of 2WikiMQA and TriviaQA. Additional training details are provided in Appendix A.6. We compare our method with 5 link tokens to the best-performing baseline, BlockAttention, under KV cache compression on QA tasks.
The evaluation results are shown in Table 3. We highlight the following findings. First, compared to LLMLingua, the modified AnLLMs compression retains substantially more information, resulting in consistently stronger performance across all QA benchmarks. Second, adding link tokens consistently mitigates the accuracy drop introduced by compression, restoring the model performance when using both KV cache compression techniques.
Table 3: Evaluation of KVLink under different compression regimes. We report the QA accuracy when using compressed cache. The value to the left of the slash (/) corresponds to a 75% compression rate and the value to the right corresponds to 50% compression.
NQ 2WikiMQA TriviaQA HotpotQA MuSiQue
75% / 50% 75% / 50% 75% / 50% 75% / 50% 75% / 50%
Prompt Compression BlockAttention 33.9 / 37.8 47.3 / 58.2 60.1 / 64.8 36.1 / 41.1 7.1 / 11.1
KVLink 35.5 / 41.6 47.8 / 58.2 61.0 / 65.7 37.3 / 46.6 6.5 / 11.5
Our Method BlockAttention 37.5 / 40.5 68.4 / 70.0 67.0 / 68.8 51.1 / 54.3 14.0 / 16.7
KVLink 40.9 / 43.0 69.9 / 69.4 68.2 / 69.3 52.4 / 55.4 14.8 / 17.3
@@ 4 Related Work
Efficient inference for LLMs. As model sizes grow, serving LLMs efficiently becomes increasingly challenging. Previous work has tackled this via model pruning [35, 36, 37, 38], quantization [39, 40, 41, 42, 43, 44], and optimized decoding algorithms such as non-autoregressive decoding [45], speculative decoding [46, 47, 48], or early-exiting [49, 50]. Like these approaches, our method also aims to enhance LLM inference efficiency by reducing inference-time computation.
KV cache reuse. Recent work has explored reusing precomputed KV caches across multiple queries, often focusing on prefix-only reuse [51, 52, 53]. A few methods extend reuse to arbitrary positions, aiming for greater efficiency. PromptCache[9] allows KV cache reuse with discontinuous position IDs but ignores cross-chunk attention, causing performance drops. It also relies on Prompt Markup Language, limiting flexibility across tasks. CacheBlend reduces positional mismatches by selectively recomputing caches for tokens with the largest value-state variance; however, it suffers from performance degradation and the non-trivial cost of recomputing 10%–20% of tokens (plus all tokens at the first layer). BlockAttention[7] removes positional embeddings by re-encoding KV caches and directly fine-tunes on block-attention data, but lacks strategies to better connect reused contexts or refine data processing, leaving performance below ideal levels.
KV compression. While our method focuses on accelerating inference by reusing KV caches, a complementary line of work aims to reduce memory overhead through cache eviction and quantization. For instance,  [54, 55, 56, 57] introduces an eviction policy that significantly lowers the memory footprint during generation, while  [58, 42, 39, 41, 59, 60, 61] investigate quantizing KV caches to minimize storage requirements with only minor performance loss. Notably, our approach can be seamlessly integrated with these state-of-the-art eviction and quantization strategies, thereby addressing both speed and memory efficiency.
Retrieve-augmented generation. Retrieval-augmented generation (RAG) is widely used to enhance the generation capabilities of language models by retrieving supporting documents. Retrievers such as dense retrievers or BM25 [22, 62, 63] are typically employed to find the most relevant documents for a given user query or context. In addition to simply concatenating the retrieved text with the input [64], there are also other methods for integrating the retrieved knowledge into the models. For instance, LongMem [65] encodes documents into caches and uses a trained side net to incorporate these knowledge caches into the context for long-context tasks. With the advances of LLMs across various NLP tasks, certain studies have focused on improving RAG with LLM [66, 67]. Some research also explores how to enhance the robustness of LLMs when processing retrieved knowledge. For example, RAFT [68] provides a fine-tuning strategy that strengthens LLMs against noisy contexts in domain-specific RAG.
@@ 5 Conclusion
In this paper, we introduced a method to improve LLM efficiency by reusing pre-computed KV caches. With precomputed KV caches for retrieved documents or context segments, our method can avoid redundant computation for overlapping contexts across different queries. In the future, we will refine our data mixture and investigate optimal fine-tuning strategies to further enhance performance. Additionally, we plan to explore real-world deployment scenarios to maximize the inference efficiency of LLMs. The limitation and societal impact sections are included in the Appendix.
@@ Appendix A Appendix
A.1 Data Mixture
We fine-tuned our model using a mixture of different datasets.
Retrieval-augmented QA.
Since RAG is the most directly relevant task, we first consider fine-tuning the model on QA tasks with retrieved documents that are separately encoded to ensure it can produce high-quality responses under this setting. More specifically, we sample questions from TriviaQA [19] and 2WikiMQA [18]. For each question, we retrieve ten relevant Wikipedia passages using Contriever [22], each encoded separately. To obtain high-quality supervision, we prompt GPT-4 to generate reference answers, which serve as the ground truth for training.
Multi-turn conversation.
Many existing instruction-following datasets [69] contain multi-turn conversations between the user and the assistant, which provides a natural way to train the model on disjointed contexts. We randomly convert earlier conversation turns into independently encoded KV caches, and the LLM is then trained to produce the appropriate responses in subsequent turns. This ensures the model to learn to integrate segmented contexts while maintaining its instruction-following capability.
Summarization.
Summarization serves as another useful setting where the model must aggregate information from independently encoded document chunks to generate a coherent summary. We adopt the XSum [70] dataset. Each document is randomly split into multiple consecutive segments, and each segment is independently encoded into KV cache. The model is trained to generate the ground-truth summary, ensuring it learns to integrate information across separate segments.
Additionally, to better preserve the model’s original capabilities, we also include a standard version of TriviaQA and 2WikiMQA, where the retrieved documents are encoded as a whole rather than separately. In this setting, the model is trained to generate the ground-truth answer conditioned on the fully concatenated context. Lastly, we incorporate the T"ulu 3 dataset [17] to preserve the instruction-following ability and a small portion of pre-training data from Fineweb [16] to preserve the language modeling capability.
@@ A.2 Implementation Details
we propose fine-tuning the LLM with a mixed dataset drawn from different sources, enabling the model to integrate disjoint context segments while preserving its original capabilities.
Figure 5 illustrates the system prompts we use for each task category during training. These prompts are generated by GPT-4 [71], and are randomly picked during training. Additionally, for the QA training data, we also shuffle the retrieved reused documents in the context A.5.
@@
For QA evaluation with separately encoded KV cache, we adopt the accuracy as the evaluation metric, following [7, 21, 72, 73], which considers the prediction is correct if any sub-string of the prediction exactly matches any gold answer.
For the evaluation of general capability preservation, we use LM Evaluation Harness [74]. We use few-shot examples for GSM8K and MMLU(8-shot for GSM8K and 5-shot for MMLU).
@@ A.3 More Implementation Details of Baselines
For BlockAttention, we process the training data with context reuse in 4 by separately encoding each reused document or context segment and concatenating the caches directly without inserting any special tokens in between. For the data with no context reuse, the data processing is the same as KVLink.
@@ A.4 Ablation Studies on Data Mixure
Table 5: Performance on different training data mixture.
@@ A.5 Impact of Answer Document Position
Empirically, we have observed that directly constructing the QA training data using Contriever [22] retrieved relevant documents is not ideal because the retriever typically places the answer-containing document near the front based on the relevance score. As a result, the fine-tuned model “learns” to find answers at the beginning of the context rather than reasoning across the entire reused context.
@@
To validate our statement, we fine-tune the Llama-3.2-3B base model on the same training data without shuffling as BlockAttention and evaluate under the [21] setup. In [21], the test set for NaturalQuestions is also built with Contriever-retrieved contexts but places the document containing the correct answer at varying positions. As shown in Figure 6, the model’s performance drops significantly whenever the ground-truth document is located further back in the context.
A.6 Training of Modified AnLLM Compression
Our training of modified AnLLM compression contains two stages: continuous pre-training and QA tuning.
Figure 7: The attention during training compression tokens.
Continuous pre-training.
This stage involves training the compression tokens to capture information from a longer preceding sequence. As shown in Figure 7, given a text sample, we first divide it into two halves. We compress the former half and calculate the cross-entropy loss on the latter half. During compression, we segment the former half into multiple text segments based on a fixed token length of 100 tokens. After each segment, we append several compression tokens. These compression tokens attend to all the preceding compression tokens and the original tokens in their segment, which distinguishes our approach from AnLLM. The original tokens in each segment maintain local attention, and the tokens in the latter half cannot attend to these original tokens, but only to the compression tokens.
QA tuning.
The goal of QA tuning is to adapt the compression tokens to compress the document and further train the model to perform QA tasks based on the compressed information. The attention mechanism during QA tuning is similar to that in continuous pre-training. The only difference is that the compression tokens of each document cannot attend to one another.
A.7 Limitations
Although our approach achieves state-of-the-art performance while improving the inference efficiency, it still has several limitations. Although KVLink can restore performance with compressed cache, there is still some performance degradation after compression. One possible solution is to conduct larger scale training for better compression and linking. Second, the size of KV cache varies a lot, because it is partially based on the document length. Therefore, it will be hard to store these cache in the vector database. Efficiently organizing and storing the pre-computed KV cache remains a challenge for large-scale deployment.
A.8 Societal Impact
A.9 License
@@
Various datasets are included for training and evaluation, and their licenses are listed below. The IFEval, NaturalQuestions, 2WikiMQA, TriviaQA, HotpotQA dataset is released under the Apache License 2.0. HumanEval, MMLU, GSM8K, HellaSwag, WinoGrande, XSum and LM-Evaluation-Harness are under the MIT License. The ARC dataset is provided under the Creative Commons Attribution Share Alike 4.0 license, and SciQ is under the Creative Commons Attribution Non-Commercial 3.0 license. The DaringAnteater dataset is under the license of Creative Commons Attribution 4.0. The MuSiQue dataset is under the Creative Commons Attribution 4.0 International license. The Samsum dataset is under the Creative Commons Attribution Non Commercial No Derivatives 4.0 license. The MultiNews dataset is under a legal agreement with LILY LAB. PiQA is licensed under the Academic Free License v. 3.0. The Tulu3-sft-mixture and fineweb datasets are under the Open Data Commons License Attribution family license. All datasets and software packages in this study are used strictly for their intended purpose—namely, training and evaluating LLMs. We confirm that no personal identifying information or offensive content appears in our materials.
A.10 LLM Usage
@@
We employed GPT-4 to assist with proofreading and improving clarity throughout the text. Nevertheless, the research concepts, analysis, and original writing remain fully authored by us.